Last year at this time, a mutual fund had a NAV of $13.20 per share. Over...

  1. Last year at this time, a mutual fund had a NAV of $13.20 per share. Over the past year the fund paid dividends of $0.70 per share and had a capital gains distribution of $1.20 per share. What is the holding period return assuming that the current NAV is $14.42?

  • 23.6

  • 30

  1. What is the fundamental value of a call with a strike price of $30 and a market price of $33?

  • $300

  1. Grant purchased one call on XYZ stock at an exercise price of $25. The market price of XYZ stock when Grant purchased the call was $24 a share. XYZ is currently priced at $30 a share. Grant paid $120 to buy the call. How much profit will Grant make if he exercises the option today and then sells the shares? Ignore all transaction-related costs

  • $38

  • $25

  1. The price of ABC stock is currently $42 per share, but in six months you expect it to rise to $50. ABC does not pay a dividend. You buy a six-month call on ABC, with a strike price of $45. The option cost $200. What holding period return do you expect on this call? Ignore transaction costs and taxes

  • 150

  • 245

please show work
Answer #1

Answer 1

holding period return = (income + (end of period value - initial value))/initial value

income = dividends + capital gains distribution= 0.7 + 1.2

initial value = 13.2, end of period value = 14.42

HPR = (0.7+1.2 + (14.42-13.2))/13.2 = 23.6

**As per HOMEWORKLIB POLICY we are entitled to attempt only one question which is the first question or a particular question number specified. Please re-post the remaining questions.